Irwin, Inc., constructed a machine at a total cost of $35 million. Construction was completed at the end of 2014 and the machine was placed in service at the beginning of 2015. The machine was being depreciated over a 10-year life using the straight-line method. The residual value is expected to be $2 million. At the beginning of 2018, Irwin decided to change to the sum-of-the-years'-digits method. Ignoring income taxes, prepare the journal entry relating to the machine for 2018. (If no entry is required for a transaction/event, select "No journal entry required" in the first account field. Enter your answers in millions rounded to 2 decimal places (i.e., 5,500,000 should be entered as 5.50).) help me because 2.4 it is wrong answer
Answer #1

Straight line depreciation per year = (35-2/10) = 3.3 per year

Accumulated depreciation (2015 to 2017) = 3.3*3 = 9.9 million

Sum of year digit = 7+6+5+4+3+2+1 = 28

Depreciation expense for 2018 = (35-2-9.9)*7/28 = 5.78 million

Date account and explanation debit credit
2018 Depreciation expense 5.78
Accumulated depreciation-Machine 5.78
(To record Dep)